# PiFlow: Principle-aware Scientific Discovery with Multi-Agent Collaboration

### v2

@@ Abstract
Large Language Model (LLM)-based multi-agent systems (MAS) demonstrate remarkable potential for scientific discovery. Existing approaches, however, often automate scientific discovery using predefined workflows that lack rationality constraints. This often leads to aimless hypothesizing and a failure to consistently link hypotheses with evidence, thereby hindering the systematic reduction of uncertainty. Overcoming these limitations fundamentally requires a principled approach to exploration. We introduce PiFlow, an information-theoretical framework, treating automated scientific discovery as a structured uncertainty reduction problem guided by principles (e.g., scientific laws). In evaluations across three distinct scientific domains – discovering nanomaterial structures, bio-molecules, and superconductor candidates with targeted properties – our method significantly improves discovery efficiency, reflected by a 73.55% increase in the Area Under the Curve (AUC) of property values versus exploration steps, and enhances solution quality by 94.06% compared to a vanilla agent system. Overall, PiFlow serves as a Plug-and-Play method, establishing a novel paradigm shift in highly efficient automated scientific discovery, paving the way for more robust and accelerated AI-driven research.
@@ 1 Introduction
Large Language Model (LLM)-based Multi-Agent Systems (MAS) have significantly impacted automated scientific discovery (Minaee et al., 2024; Wang et al., 2023; Zhang et al., 2024c) across a wide range of fundamental fields, including chemistry (Liu et al., 2024; Ghafarollahi & Buehler, 2024a; Yang et al., 2024c; Inoue et al., 2024), biology (Xiao et al., 2024; Nagarajan et al., 2025; Averly et al., 2025; Ghafarollahi & Buehler, 2024b), physics (Jaiswal et al., 2024), and material science (Takahara et al., 2025; Ghafarollahi & Buehler, 2025; Ansari et al., 2024; Wan et al., 2024; Zhang et al., 2024b).
Although proficient in executing experiments within predefined workflows (Lu et al., 2024; Lai & Pu, 2025), these systems often generate hypotheses that lack clear direction, leading to the uncertainty establishing clear links between hypotheses and their supporting or refuting evidence (AI4Science & Quantum, 2023; Zhou et al., 2024; Baek et al., 2024; Schmidgall et al., 2025; Prabhakar et al., 2025; Xie et al., 2023a). Such a disconnect indicates inefficient exploration. Moreover, many of these approaches are tailored for specific tasks, often relying on meticulous prompt engineering that heavily incorporates domain knowledge (as detailed in Appendix D). Consequently, their ability to adapt to new scientific domains is often limited without substantial modifications (Zhang et al., 2025; Kumbhar et al., 2025). These issues culminate in three primary challenges: (a) aimless hypothesizing; (b) unmaintained connections between hypotheses and evidence during exploration, hindering systematic validation and (c) limited generalization ability, where systems effective in one scenario (e.g., material science) often require substantial rework to be applicable in others.
@@
This principle-aware approach yields systematic information gain: PiFlow selects high-potential principles and then guides hypothesizing via three actions, i.e., exploring, validating, or refining their scope and formulation. Thus, PiFlow progressively optimizes its guiding scientific principles to effectively steer hypothesizing. Furthermore, leveraging its Min-Max optimization, PiFlow theoretically achieves cumulative regret growth of $\mathcal{O}(\sqrt{T})$ over $T$ exploration steps (a detailed proof is provided in Appendix F). This sublinear regret underscores its guaranteed efficiency in navigating complex discovery landscapes. In summary, our contributions are:
@@
    We develop PiFlow, a Plug-and-Play framework that seamlessly integrates with existing MAS to enable focused exploration, thereby enhancing discovery efficiency and flexibility.
@@
    We conduct extensive experiments across three distinct scenarios, demonstrating the broad applicability and significant performance improvements achieved by PiFlow.
@@ 2.1 Language Models for Scientific Discovery
Recently, large language models (LLM) have advanced scientific discovery with automation and rational design (Ren et al., 2025; Ma et al., 2024). The internal knowledge of LLMs has demonstrated promising capability in focused chemical and material discovery (Yang et al., 2024c; Zhou et al., 2024; Pu et al., 2024; Ghafarollahi & Buehler, 2024c). While tool-integrated LLMs like SciAgents (Ghafarollahi & Buehler, 2024d), DARWIN (Xie et al., 2023a) and HoneyComb (Zhang et al., 2024a) improve domain-specific reasoning and recall of factual insights, they still struggle to integrate physicochemical laws effectively when refining insights for design, risking biased proposals and inefficient exploration due to inherent hallucinations of LLMs (Zhang et al., 2023b). Human-AI frameworks address this issue by leveraging the knowledge of domain experts (Reddy & Shojaee, 2024; Eythorsson & Clark, 2025), yet remain limited to the scope of hypothesis generation (Alkan et al., 2025), leading to insufficient exploration of complex chemical spaces (Luo et al., 2025). Surveys highlight persistent gaps in efficiency and interpretability (Zhang et al., 2024c; Han et al., 2024), underscoring the need for principled scientific discovery management beyond automated LLM reasoning (Ramos et al., 2024).
@@ 2.2 Approaches of Multi-Agent Collaboration
Multi-agent systems (MAS) show promise for complex tasks (Lu et al., 2024; Ghafarollahi & Buehler, 2024d; Ni & Buehler, 2023), yet their application to scientific discovery reveals limitations in current collaboration mechanisms (Tran et al., 2025). Rule-based methods (Zhang et al., 2023a) offer consistency but their predefined rules lack the flexibility to incorporate nuanced scientific principles (e.g., physicochemical laws) or adapt to unexpected findings, hindering dynamic exploration. Role-playing approaches (Tran et al., 2025; He et al., 2024) leverage agent expertise, yet rigid roles can impede adaptation in scientific research (Ramirez-Medina et al., 2025; Lu et al., 2024; Ghafarollahi & Buehler, 2024d), and ensuring collective adherence to scientific principles when interpreting experimental insights is challenging. Model-based methods (Xu et al., 2023; Mu et al., 2023; Li et al., 2023) aim for adaptability by learning from uncertainty, but struggle to build world models that accurately capture complex scientific phenomena and integrate guiding laws (Hao et al., 2023), thereby impairing the balance between information perception and strategic, principle-guided reasoning.
Consequently, existing MAS paradigms often lack a dedicated awareness and systematic application of scientific principles during hypothesis generation and refinement (Gridach et al., 2025; Luo et al., 2025; Reddy & Shojaee, 2024; Su et al., 2024). This highlights a critical need for an explicitly principle-aware multi-agent collaboration framework. Our work addresses this gap, proposing a method where the collaborative discovery process is robustly guided by scientific principles to achieve more efficient and reliable outcomes.
@@ 3.2 Architecture
    Hypothesis. Initially, a set of dynamically growing candidate principles $\mathcal{P}=\{p_{1},p_{2},p_{3}\}$ (see Definition 3.1) potentially proposed by domain experts (see Appendix P) or extracted by LLMs, is established. In each iteration $t$, $\mathcal{A}_{H}$ proposes a testable hypothesis, $h_{t}$, grounded in a selected principle $p_{i}\in\mathcal{P}$. The proposal is supported by structured rationales (comprising major and minor premises, following Eisape et al. (2023), see Appendix T.2).
@@
This closed-loop feedback mechanism ensures that the Hypothesis-Validation cycle is continuously and adaptively steered by strategic insights derived from the system’s cumulative experience, as embodied in $\mathcal{T}_{t}$. We provide a detailed analysis of the distinction from prompt engineering at Appendix E, and an illustrative example at Appendix H.
Plug-and-Play Modularity of PiFlow. The hypothesis steering mechanism above, which includes PiFlow and its interfacing Planner agent $\mathcal{A}_{P}$, has been intentionally engineered as a modular, Plug-and-Play system. This architectural choice ensures that principle-aware guidance can be readily integrated to enhance MAS engaged in scientific discovery as a part of prompts. As demonstrated in Appendix K, we successfully integrate PiFlow with existing ChemToolAgent (Yu et al., 2024) without any architecture modifications. This adaptability positions PiFlow as a pivotal enhancement for automated scientific inquiry.
@@ 3.3 Min-Max Optimization in PiFlow
Minimizing cumulative fegret (exploitation). The first term, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents the cumulative regret over $T$ iterations. Here, $v^{*}$ is a theoretical optimal outcome value, and $f^{*}(h_{t})$ is the outcome from hypothesis $h_{t}$. By minimizing this term, the policy $\pi$ is driven to exploit known high-potential principles and hypotheses to achieve outcomes as close as possible to the optimum $v^{*}$. This encourages the refinement and validation of promising avenues.
@@
Building upon the theoretical framework above, we derive a computationally tractable algorithm (Algorithm 1) that serves as a principled approximation of the abstract Min-Max objective in Eq. 1. The full rationale and derivation for this approximation are detailed in Appendix G.
In summary, the Min-Max adversarial formulation underpinning PiFlow provides strong theoretical guarantees, notably sublinear regret bounds (formalized in Theorem 1, with proof in Appendix F). Importantly, its operational behavior, which involves practical approximations of this Min-Max solution, demonstrates consistent alignment with these theoretical expectations, as empirically validated in Section 5.4.
Theorem 1 (Informal).
@@ 4.1 Settings
To ensure a fair comparison and focus on core capabilities, all agents utilize QwenMax (Yang et al., 2024a) as the base LLM and are prohibited from accessing external search tools. The complete experimental setup is detailed in Appendix S, and the role-playing prompts for PiFlow are provided in Appendix T.
@@ 4.2 Experimental Scenarios
To comprehensively assess PiFlow’s performance, we design three scenarios that represent canonical challenges in scientific exploration: optimization in continuous, discrete, and mixed parameter spaces:
Nanohelix Optimization (NHO). We use a surrogate model ($r^{2}=0.98$) trained on DFT-simulated data following Wu et al. (2025) to predict nanohelix chirality from four continuous geometric parameters, enabling efficient exploration of the design space (Appendix S.1).
Molecular Bio-activity Optimization (MBO). We build a surrogate model ($r^{2}=0.91$) to predict bio-activity from SMILES strings, trained on 50,000 molecules from ChEMBL35 (Zdrazil et al., 2023). This facilitates high-throughput screening in a discrete chemical space. (Appendix S.2).
Superconductor Optimization (SPO). Following  Hamidieh (2018), we train a surrogate model ($r^{2}=0.91$) to map a material’s mixed continuous and discrete compositional features to its critical temperature ($T_{c}$), accelerating the discovery of room-temperature superconductors. (Appendix S.3).
@@ 4.3 Baselines
To evaluate the strategic guidance of PiFlow under uncertainty, we therefore benchmark against the following baselines: (1) Reasoning and Acting (ReAct) (Yao et al., 2022). ReAct enables agents to iteratively formulate hypotheses, design/execute experiments, and interpret results. It represents a foundational approach to structured reasoning. (2) Meta Plan Optimization (MPO) (Xiong et al., 2025). MPO employs a trained LLM planner that provides high-level, general guidance. This serves as a direct counterpoint to PiFlow’s explicit, structured mechanism for principled uncertainty reduction. (3) Vanilla Agent System (Vanilla). This baseline consists of an MAS operating without any principled strategic oversight. It is intended to establish a performance floor, demonstrating the limitations of unguided exploration reliant solely on agent role-playing.
While other advanced frameworks exist, such as Agent-Oriented Planning (AOP) (Li et al., 2024) and Reason for Future, Act for Now (RAFA) (Liu et al., 2023), their design objectives diverge from the scope of our evaluation. AOP is optimized for tasks with well-defined structures rather than initial epistemic uncertainty, while RAFA focuses on completing predefined goals rather than the iterative, de novo evidence-gathering process central to PiFlow. Therefore, our curated selection of baselines is specifically designed to isolate and evaluate our contribution to discovery efficiency in uncertain environments.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Large Language Model (LLM)-based Multi-Agent Systems (MAS) have significantly impacted automated scientific discovery (Minaee et al., 2024; Wang et al., 2023; Zhang et al., 2024c) across a wide range of fundamental fields, including chemistry (Liu et al., 2024; Ghafarollahi and Buehler, 2024a; Yang et al., 2024b; Inoue et al., 2024), biology (Xiao et al., 2024; Nagarajan et al., 2025; Averly et al., 2025; Ghafarollahi and Buehler, 2024b), physics (Jaiswal et al., 2024), and material science (Takahara et al., 2025; Ghafarollahi and Buehler, 2025; Ansari et al., 2024; Wan et al., 2024; Zhang et al., 2024b).
Although proficient in executing experiments within predefined workflows (Lu et al., 2024; Lai and Pu, 2025), these systems often generate hypotheses that lack clear direction, leading to the uncertainty in establishing clear links between hypotheses and their supporting or refuting evidence (AI4Science and Quantum, 2023; Zhou et al., 2024; Baek et al., 2024; Schmidgall et al., 2025; Prabhakar et al., 2025; Xie et al., 2023a). Such a disconnect indicates inefficient exploration. Moreover, many of these approaches are tailored for specific tasks, often relying on meticulous prompt engineering that heavily incorporates domain knowledge (as detailed in Appendix A). Consequently, their ability to adapt to new scientific domains is often limited without substantial modifications (Zhang et al., 2025; Kumbhar et al., 2025). These issues culminate in three primary challenges: (a) aimless hypothesizing; (b) unmaintained connections between hypotheses and evidence during exploration, hindering systematic validation and (c) limited generalization ability, where systems effective in one scientific scenario (e.g., material science) often require substantial rework to be applicable in others.
@@
By operationalizing the Min-Max optimization, PiFlow systematically steers discovery via three strategic actions: explore, validate, or refine, thereby maximizing information gain while minimizing cumulative regret. This mechanism theoretically guarantees a sublinear regret bound of $\mathcal{O}(\sqrt{T})$ (proof in Appendix C), ensuring that the system avoids aimless exploration and efficiently converges to optimal solutions even in complex, high-dimensional landscapes.
[…]
In summary, our contributions are:
@@
    We develop PiFlow, a Plug-and-Play framework that seamlessly integrates with existing MAS to enable focused hypothesis exploration, thereby enhancing discovery efficiency and flexibility.
@@
    We conduct extensive experiments across three distinct scientific discovery scenarios, demonstrating the broad applicability and significant performance improvements achieved by PiFlow.
@@ 2.1 Language Models for Scientific Discovery
Recently, large language models (LLM) have advanced scientific discovery with automation and rational design (Ren et al., 2025; Ma et al., 2024). The internal knowledge of LLMs has demonstrated promising capability in focused chemical and material discovery (Yang et al., 2024b; Zhou et al., 2024; Pu et al., 2024; Ghafarollahi and Buehler, 2024c). While tool-integrated LLMs like SciAgents (Ghafarollahi and Buehler, 2024d), DARWIN (Xie et al., 2023a) and HoneyComb (Zhang et al., 2024a) improve domain-specific reasoning and recall of factual insights, they still struggle to integrate physicochemical laws effectively when refining insights for design, risking biased proposals and inefficient exploration due to inherent hallucinations of LLMs (Zhang et al., 2023b). Human-AI frameworks address this issue by leveraging the knowledge of domain experts (Reddy and Shojaee, 2024; Eythorsson and Clark, 2025), yet remain limited to the scope of hypothesis generation (Alkan et al., 2025), leading to insufficient exploration of complex chemical spaces (Luo et al., 2025). Surveys highlight persistent gaps in efficiency and interpretability (Zhang et al., 2024c; Han et al., 2024), underscoring the need for principled scientific discovery management beyond automated LLM reasoning (Ramos et al., 2024).
@@ 2.2 Approaches of Multi-Agent Collaboration
Multi-agent systems (MAS) show promise for complex tasks (Lu et al., 2024; Ghafarollahi and Buehler, 2024d; Ni and Buehler, 2023), yet their application to scientific discovery reveals limitations in current collaboration mechanisms (Tran et al., 2025). Rule-based methods (Zhang et al., 2023a) offer consistency but their predefined rules lack the flexibility to incorporate nuanced scientific principles (e.g., physicochemical laws) or adapt to unexpected findings, hindering dynamic exploration. The method of AI Researcher (Tang et al., 2025), and role-playing approaches (Tran et al., 2025; He et al., 2024) leverage agent expertise, yet rigid roles can impede adaptation in scientific research (Ramirez-Medina et al., 2025; Lu et al., 2024; Ghafarollahi and Buehler, 2024d), and ensuring collective adherence to scientific principles when interpreting experimental insights is challenging. Model-based methods (Xu et al., 2023; Mu et al., 2023; Li et al., 2023) aim for adaptability by learning from uncertainty, but struggle to build world models that accurately capture complex scientific phenomena and integrate guiding laws (Hao et al., 2023), thereby impairing the balance between information perception and strategic, principle-guided reasoning.
Consequently, existing MAS paradigms often lack a dedicated awareness and systematic application of scientific principles during hypothesis generation and refinement (Gridach et al., 2025; Luo et al., 2025; Reddy and Shojaee, 2024; Su et al., 2024). This highlights a critical need for an explicitly principle-aware multi-agent collaboration framework. Our work addresses this gap, proposing a method where the collaborative discovery process is robustly guided by scientific principles to achieve more efficient and reliable outcomes.
@@ 3.2 Architecture
    Hypothesis. Initially, a set of dynamically growing candidate principles $\mathcal{P}=\{p_{1},p_{2},p_{3}\}$ (see Definition 3.1) potentially proposed by domain experts (see Appendix M) or extracted by LLMs, is established. In each iteration $t$, $\mathcal{A}_{H}$ proposes a testable hypothesis, $h_{t}$, grounded in a selected principle $p_{i}\in\mathcal{P}$. The proposal is supported by structured rationales (comprising major and minor premises, following Eisape et al. (2023), see Appendix Q.2).
@@
This closed-loop feedback mechanism ensures that the Hypothesis-Validation cycle is continuously and adaptively steered by strategic insights derived from the system’s cumulative experience, as embodied in $\mathcal{T}_{t}$. We provide a detailed analysis of the distinction from prompt engineering at Appendix B, and an illustrative example at Appendix E.
Plug-and-Play Modularity of PiFlow. The hypothesis steering mechanism above, which includes PiFlow and its interfacing Planner agent $\mathcal{A}_{P}$, has been intentionally engineered as a modular, Plug-and-Play system. This architectural choice ensures that principle-aware guidance can be readily integrated to enhance MAS engaged in scientific discovery as a part of prompts. As demonstrated in Appendix H, we successfully integrate PiFlow with existing ChemToolAgent (Yu et al., 2024) without any architecture modifications. This adaptability positions PiFlow as a pivotal enhancement for automated scientific inquiry.
@@ 3.3 Min-Max Optimization in PiFlow
Minimizing cumulative regret (exploitation). The first term, $\sum_{t=1}^{T}(v^{*}-f^{*}(h_{t}))$, represents the cumulative regret over $T$ iterations. Here, $v^{*}$ is a theoretical optimal outcome value, and $f^{*}(h_{t})$ is the outcome from hypothesis $h_{t}$. By minimizing this term, the policy $\pi$ is driven to exploit known high-potential principles and hypotheses to achieve outcomes as close as possible to the optimum $v^{*}$. This encourages the refinement and validation of promising avenues.
@@
Building upon the theoretical framework above, we derive a computationally tractable algorithm (Algorithm 1). Since precise computation of information gain (Eq. 1) is intractable, we employ Semantic Embedding Distance as a computationally efficient proxy, justified in Appendix D.
In summary, the Min-Max adversarial formulation underpinning PiFlow provides strong theoretical guarantees, notably sublinear regret bounds (formalized in Theorem 3.2, with proof in Appendix C). Importantly, its operational behavior, which involves practical approximations of this Min-Max solution, demonstrates consistent alignment with these theoretical expectations, as empirically validated in Section 5.4.
Theorem 3.2 (Informal).
@@ 4.1 Settings
To ensure a fair comparison and focus on core capabilities, all agents utilize QwenMax (Yang et al., 2024a) as the base LLM and are prohibited from accessing external search tools. The complete experimental setup is detailed in Appendix P, and the prompts are provided in Appendix Q.
@@ 4.2 Experimental Scenarios
To comprehensively assess the performance, we design three scenarios that represent canonical challenges in scientific exploration: optimization in continuous, discrete, and hybrid parameter spaces, as detailed below.
Nanohelix Optimization (NHO). We use a surrogate model ($r^{2}=0.98$) trained on DFT-simulated data following Wu et al. (2025) to predict chirality of a nanohelix material from four continuous geometric parameters (see Appendix P.1).
Molecular Bio-activity Optimization (MBO). We use a surrogate model ($r^{2}=0.91$) trained on 50,000 molecules from ChEMBL35 (Zdrazil et al., 2023) database to predict bio-activity of a molecule, quantified by pChEMBL value, in a discrete chemical space (see Appendix P.2).
Superconductor Optimization (SPO). Following  Hamidieh (2018), we train a surrogate model ($r^{2}=0.91$) to map a material’s hybrid compositional features, e.g., Bi2Sr2Ca4Cu8O19, to its critical temperature ($T_{c}$). As reported (Hamidieh, 2018), this task is aiming for accelerating the discovery of room-temperature superconductors (see Appendix P.3 for more details).
Table 1: Performance Comparison on Four Scientific Discovery Tasks. Performance is measured by AUC (Area Under Curve, %) and SQ (Solution Quality, %) and mean pmpm std is reported. Bold values denote the best result in each column, underlined values denote the second-best result. The rightmost columns show task-averaged performance, with red percentages indicating the relative performance gap of baseline methods compared to our PiFlow.
Method NHO (Continuous) MBO (Discrete) SPO (Hybrid) Average
AUC (%) $\uparrow$ SQ (%) $\uparrow$ AUC (%) $\uparrow$ SQ (%) $\uparrow$ AUC(%) $\uparrow$ SQ (%) $\uparrow$ AUC (%) SQ (%)
ReAct 35.85 $\pm$7.12 41.96 $\pm$0.82 29.61 $\pm$9.74 43.11 $\pm$9.98 5.29 $\pm$0.69 6.41 $\pm$0.96 23.58 $\downarrow$ 52.70% 30.49 $\downarrow$ 53.38%
Vanilla 35.96 $\pm$22.38 46.76 $\pm$7.29 29.71 $\pm$12.66 49.22 $\pm$8.30 11.39 $\pm$11.33 14.16 $\pm$13.37 25.69 $\downarrow$ 48.49% 36.71 $\downarrow$ 43.87%
MPO 43.99 $\pm$2.79 51.29 $\pm$7.77 31.18 $\pm$9.12 57.28 $\pm$5.85 12.68 $\pm$7.75 33.20 $\pm$23.75 29.28 $\downarrow$ 41.27% 47.26 $\downarrow$ 27.75%
AI Researcher 46.45 $\pm$1.98 53.12 $\pm$0.87 42.72 $\pm$12.71 95.66 $\pm$6.08 16.36 $\pm$2.22 25.69 $\pm$2.64 35.18 $\downarrow$ 29.45% 58.16 $\downarrow$ 11.08%
The AI Scientist v2 49.27 $\pm$1.51 56.67 $\pm$4.73 36.32 $\pm$8.67 62.47 $\pm$6.55 28.43 $\pm$2.55 29.85 $\pm$2.18 38.01 $\downarrow$ 23.78% 49.66 $\downarrow$ 24.07%
PiFlow (ours) 63.51 $\pm$11.18 76.82 $\pm$4.54 64.57 $\pm$23.65 84.55 $\pm$29.63 21.51 $\pm$2.80 34.85 $\pm$1.19 49.86 65.41
Figure 3: Optimization trajectories across distinct scientific domains. The curves depict the objective values over 24 budget steps. Methods like The AI Scientist v2 (Yamada et al., 2025) remains stuck (e.g. from “Bi2Sr2Ca4Cu8O18” to “Bi2Sr2Ca4Cu8O19” in SPO). However, PiFlow exhibits characteristic oscillatory patterns indicative of its active principle-aware exploration strategy, attains the best performance of 65.41% SQ and 49.86% AUC on average.
@@ 4.3 Baselines
To evaluate the strategic guidance of PiFlow under uncertainty, we therefore benchmark against the following baselines: (1) Reasoning and Acting (ReAct) (Yao et al., 2022). ReAct enables agents to iteratively formulate hypotheses, design/execute experiments, and interpret results. It represents a foundational approach to structured reasoning. (2) Meta Plan Optimization (MPO) (Xiong et al., 2025). MPO employs a trained LLM planner that provides high-level, general guidance. This serves as a direct counterpoint to PiFlow’s explicit, structured mechanism for principled uncertainty reduction. (3) Vanilla Agent System (Vanilla). This baseline consists of an MAS operating without any strategic oversight. It is intended to establish a performance floor, demonstrating the limitations of unguided exploration reliant solely on agent role-playing. (4) AI Researcher (Tang et al., 2025), a multi-agent research system for autonomous scientific innovation. (5) The AI Scientist v2 (Yamada et al., 2025), an end-to-end agentic system that leverages an tree-search methodology for prompting research progress. We adapt these two AI Scientist methods into our tasks by excluding any information retrieval from external resources and manuscript drafting for a fair comparison.
While other advanced frameworks exist, such as Co-Scientist (Gottweis et al., 2025) and Kosmos (Mitchener et al., 2025), their objectives diverge from the scope of de novo evidence-gathering process central to PiFlow. Therefore, our selection of baselines is specifically designed to isolate and evaluate our contribution to discovery efficiency in uncertain environments.